OpenStudy (anonymous):

An arrow is shot straight up in the air at an initial speed of 15.0 m/s. After how much time is the arrow moving downward at a speed of 8.00 m/s?

OpenStudy (anonymous):

delta t=delta v/a= (vf-v0)/-g (-8-15)/-9.81= 2.34 seconds